# Towards Optimizing with Large Language Models

### v1

@@
Abstract
In this work, we conduct an assessment of the optimization capabilities of LLMs across various tasks and data sizes. Each of these tasks corresponds to unique optimization domains, and LLMs are required to execute these tasks with interactive prompting. That is, in each optimization step, the LLM generates new solutions from the past generated solutions with their values, and then the new solutions are evaluated and considered in the next optimization step. Additionally, we introduce three distinct metrics for a comprehensive assessment of task performance from various perspectives. These metrics offer the advantage of being applicable for evaluating LLM performance across a broad spectrum of optimization tasks and are less sensitive to variations in test samples. By applying these metrics, we observe that LLMs exhibit strong optimization capabilities when dealing with small-sized samples. However, their performance is significantly influenced by factors like data size and values, underscoring the importance of further research in the domain of optimization tasks for LLMs.
I Introduction
@@
In recent years, Large Language Models have demonstrated exceptional capabilities in reasoning across a variety of natural language-based tasks  [5]. However, their potential extends beyond multiple-choice questions or single-question answering. This work explores LLMs’ effectiveness in optimization across diverse tasks and data sizes. Optimization, in this context, refers to the process of iteratively generating and evaluating solutions to improve upon a given objective function.
@@
Our research assesses how well LLMs perform in interactive optimization settings, where each optimization step involves the generation of new solutions based on past solutions and their associated values. We conduct our study with four different types of optimization algorithms: Gradient Descent, Hill Climbing, Grid Search, and Black Box Optimization. In order to provide a comprehensive evaluation of LLM performance, we introduce three distinct metrics. These metrics provide a multifaceted view of task performance and are applicable across a broad spectrum of optimization tasks, reducing sensitivity to sample variations.
Our preliminary findings indicate that LLMs exhibit impressive optimization capabilities, particularly when confronted with tasks involving small-sized data samples. However, the performance of these models is notably influenced by factors, including the sample size and the range of values involved in the process. These observations underscore the need for further research and exploration within the domain of optimization tasks tailored for Large Language Models.
To summarize, we aim to make the following contributions in this preliminary study: (1) Exploring LLMs in interactive optimization scenarios with iterative prompting. (2) Introduce three novel evaluation metrics designed to provide a comprehensive assessment of LLM performance in optimization tasks. (3) Delve into factors that influence LLM performance in optimization, with a particular emphasis on the impact of data size and task type.
II Related Work
@@
In various optimization scenarios, the utilization of Large Language Models (LLMs) has become indispensable for the development of optimization algorithms or agent systems capable of handling complex and informative text-based feedback [1, 3, 6]. In this section, we summarize three significant related works that leverage LLMs to tackle optimization and reinforcement learning challenges:
  1. 1.
@@
    Optimization by PROmpting (OPRO): OPRO harnesses LLMs as versatile optimizers by describing optimization tasks in natural language prompts. It iteratively generates and evaluates solutions from these prompts, demonstrating superior performance on tasks like linear regression and traveling salesman problems. OPRO outperforms human-designed prompts by up to 50% on challenging tasks.
  2. 2.
@@
    Reflexion: Reflexion introduces a novel framework for training language agents that rely on linguistic feedback rather than traditional reinforcement learning. It achieves remarkable results, such as a 91% pass@1 accuracy on coding tasks, surpassing previous state-of-the-art models by 11%.
  3. 3.
@@
    EvoPrompt: EvoPrompt automates prompt optimization by connecting LLMs with evolutionary algorithms. This approach outperforms human-designed prompts by up to 25% and existing automatic prompt generation methods by up to 14%. It highlights the synergy between LLMs and conventional algorithms.
These works collectively showcase the adaptability and effectiveness of LLMs in addressing optimization and learning challenges across various domains.
III Methodology
We demonstrate that LLMs with iterative prompting can serve as optimizers and could be formulated similarly to different types of optimization algorithms. In particular, we present a case study on Gradient-Descent, Hill-Climbing (Heuristic), Grid-Search, and Black-Box Optimization (LLMs as a black box optimizer) and conduct an assessment of the optimization capabilities of Large Language Models across a range of distinct parameter search tasks.
3-1 Tasks
@@
We design four optimization tasks that require the model to algorithmically search for the optimal value of parameters. These tasks encompass Gradient-Descent, Hill-Climbing, Grid-Search, and Black-Box Optimization, each representing unique optimization domains: gradient-based, meta-heuristics, decision-theoretic, and Bayesian. In terms of parameter types, Grid-Search and Hill-Climbing involve discrete search spaces, while Gradient-Descent and Black-Box Optimization tackle continuous search spaces.
@@
In the Gradient-Descent task, we instruct Large Language Models to undertake a conventional gradient descent optimization process based on the loss function they have defined. Language Models need to compute the gradient and update the parameters using the gradient information and the learning rate given. This task assesses the model’s proficiency in advanced calculations and its grasp of the principles of gradient descent.
@@
In the Hill-Climbing task, we introduce custom instructions in the search algorithm to evaluate the LLM’s capability to adhere to predefined rules they have not seen before. LLMs start with an initial solution and iteratively explore nearby solutions by making small incremental changes. In our task, neighboring solutions are generated by selecting a specific element within the solution and either increasing or decreasing it by one each time. Subsequently, the neighbor solution with the minimum loss is chosen as the new solution and passed to the next iteration.
@@
In the Grid-Search task, Large Language Models are tasked with generating all grid points and systematically searching for the point that results in the lowest loss according to the given loss function. This task assesses the LLM’s ability to conduct exhaustive searches and locate optimal solutions within a predefined search space.
@@
In the Black-Box Optimization task, we treat the LLMs as black boxes that try to fit an unknown loss function. We provide the LLM with a limited set of solutions, each paired with its respective true loss value. The LLM’s objective is to discover a new solution that has a lower loss than the existing solutions in each iteration by themselves. This task evaluates the LLM’s ability to make informed decisions and optimize in a less transparent or more abstract problem-solving context.
Figure 1: Overview of our prompting strategy. (1) Define the loss function. (2) Given instruction prompt, LLM generates solution and pass result to next iteration. (3) Evaluate LLM final output.
3-2 Prompting Strategy
@@
We applied Chain of Thoughts as our prompting method. The Language model will accomplish each step with reasoning thoughts as intermediate outputs. Each optimization iteration is composed of three steps: (1) Define the loss function. (2) Given an instruction prompt, LLM generates a solution and passes the result to the next iteration. (3) Evaluate LLM final output. In each of these tasks (optimization algorithm), LLMs are initially required to formulate a loss function based on the provided sample. Following the formulation of this loss function, they embark on a process of iterative exploration to determine the optimal values for the variables that will minimize the loss associated with the function. Figure 1 shows an overview of how LLM performs optimization in interactive settings. The following illustrates a sample meta-prompt for the prompting strategy. Detailed meta prompt instructions can be found in Appendix A. Complete prompt examples for each optimization algorithm can be found in Appendix B.
Gradient-Descent *Define Loss function User : Given the data points (y1, y2, y3) = (2,6,0), what is the MSE loss function with respect to the ŷs for a hypothetical set of predicted ŷs values? LLM : The MSE loss function for the given data points (y1, y2, y3) = (2,6,0) with respect to ŷs is: $\frac{1}{3}[(\hat{y}_{1}-2)^{2}+(\hat{y}_{2}-6)^{2}+(\hat{y}_{3}-0)^{2}]$First Iteration User : Please minimize the loss function using gradient descent with learning rate of 0.1 at point (ŷ1, ŷ2, ŷ3, …..) = (2,3,4). What is the point we eventually end up after one update? Your answer includes two parts an explanation with calculation and a short answer of result. LLM : Explanation : To minimize the Mean Squared Error (MSE) loss function using gradient descent … Short Answer: After calculation, the next update point is (ŷ$1_{new}$, ŷ$2_{new}$, ŷ$3_{new}$) = (1.92, 3.05, 3.94) *Second Iteration User : Please minimize the loss function using gradient descent with learning rate of 0.1 at point (ŷ1, ŷ2, ŷ3, …..) = (1.92, 3.05, 3.94). What is the point we eventually end up after one update? Your answer includes two parts an explanation with calculation and a short answer of result. LLM : Explanation : To minimize the Mean Squared Error (MSE) loss function using gradient descent … Short Answer: After calculation, the next update point is (ŷ$1_{new}$, ŷ$2_{new}$, ŷ$3_{new}$) = (1.81, 3.12, 3.41)
IV Experimental
4-1 Dataset
In the experiment, we generate synthetic datasets where each instance belongs to $[0,10]^{d}$. In the experiments, we create five datasets with $d$ values chosen from the set $\{3,6,12,24,48\}$ in order to investigate sensitivity to the size of input data points.
We then apply these data to a loss function (Mean Squared Error) and find the true solution for each parameter search task. These true solutions and their losses serve as the ground truth for the tasks and are used as a benchmark to compare with the solutions found by Large Language Models afterward.
4-2 Detailed Settings
In our experiment, we performed five repetitions of the test for each sample in the dataset, with the LLM conducting ten iterations of parameter search in each repetition. We excluded excessively biased results to prevent our metrics from being skewed by a minority of poorly performing test outcomes. All the samples have been confirmed to exhibit a significant reduction in their respective loss within ten iterations when optimized by the ground truth algorithm.
@@ 4-2 Detailed Settings
We assess the performance of Large Language Models: GPT-turbo-3.5, GPT-4. To create an interactive environment, we utilize the chat-mode of GPTs, where the entire conversation history serves as the prompt. This approach enables the LLMs to retain the memory of prior search results and reasoning paths from past interactions. With each iteration, new instructions are added to the ongoing conversation records and the entire dialogue is presented as the prompt. If the dialogue exceeds the token limit, we will remove the earliest portions of the conversation to ensure compliance with the limit. See Appendix B for detailed examples.
We set the model temperature to 0.8, which adds randomness to the generated responses. This intentional randomness helps us evaluate the stability of LLM solutions when provided with identical initial settings in the search. Due to resource and token length limitations, we only test the Gradient Descent task and BlackBox Optimization task on all datasets and LLMs. For Grid Search and Hill Climbing tasks, experiments are conducted only on datasets with sample sizes 3, 6, and 12 and exclusively run on GPT-turbo-3.5.
V Metric Design
To comprehensively evaluate task performance from different aspects, we designed three distinct metrics. The first is the $goal$ $metric$, gauging the LLM’s advancement in accomplishing the optimization task. The second is the $policy$ $metric$, assessing the proximity of the final model output to the ground truth. Lastly, the $uncertainty$ $metric$, which quantifies the level of variability in the LLM’s solutions when exposed to identical conditions.
These metrics have the advantage of being versatile in evaluating LLM performance across various optimization tasks, streamlining the process for practitioners to assess multiple task performances concurrently. Additionally, these metrics are less sensitive to variations in test samples because of the use of ratio measures rather than difference measures.
5-1 Goal Metric
One of our objectives in this study is to evaluate the optimization capabilities of Large Language Models. To achieve this, we employ the goal metric, which serves as a quantitative measure of the extent to which the optimization process leads to improvements in objective function values. We define the goal metric of a test sample $j$ as :
$G_{j}=\frac{1}{N}\sum_{i=1}^{N}\frac{loss_{LLM,init}-loss_{LLM,i}}{loss_{LLM,init}}$
@@ 5-1 Goal Metric
5-2 Policy Metric
Beyond self-improvement, it’s essential to evaluate the Large Language Models’ ability to function akin to a truth model algorithm. The metric designed for this evaluation is the policy metric, which serves as an indicator of the LLM’s proficiency in adhering to task-specific instructions. We define the policy metric of a test sample $j$ as :
$P_{j}=\frac{1}{N}\sum_{i=1}^{N}\frac{loss_{LLM,i}-loss_{truth}}{loss_{truth}}$
where $loss_{LLM,i}$ is the LLM output loss of trial $i$, $loss_{truth}$ is the ground truth of sample $j$ and $N$ is the number of trials. The policy metric measures the disparity between the ground truth and the LLM’s output. A lower policy metric value indicates a more effective alignment of the LLM’s actions with the prescribed guidelines. When the value is negative, it means that LLM’s performance surpasses the ground truth.
5-3 Uncertainty Metric
Stability is a crucial characteristic in optimization tasks. We hope that the LLMs produce identical results in every trial involving the same sample, even under conditions with temperatures greater than zero. We define the uncertainty metric of a test sample $j$ as :
$U_{j}=\frac{1}{N}\sum_{i=1}^{N}(loss_{LLM,i}-\overline{loss_{LLM}})^{2}$
@@ 5-3 Uncertainty Metric
VI Results and Analysis
@@
We summarize the outcomes of our experiment and subsequently examine the common trends observed across all experiments. In every plot, the x-axis displays the number of data points within each sample. In the case of the goal metric and policy metric plots, the y-axis illustrates the average metric value for the respective tasks, while the shaded area in a lighter color delineates the confidence interval of the metric, denoted as $[value-std,value+std]$. As for the uncertainty metric plot, the y-axis showcases the uncertainty metric value, which corresponds to the standard deviation of the LLM final solution loss.
6-1 Results
Figure 2 presents the performance of GPT-turbo-3.5 across all tasks. The Grid-Search task doesn’t appear in the goal metric graph due to its non-iterative nature. Similarly, Black-Box Optimization isn’t represented in the policy metric graph because ground truth is unattainable.
Figure 2: GPT-turbo-3.5 average zero-shot performance across tasks
@@ 6-1 Results
We can see that GPT-turbo-3.5 exhibits optimization capabilities in most cases. Impressively, in the Gradient-Descent task, GPT-turbo-3.5 even surpasses the ground truth, particularly in the case of the 6 data points sample. It’s also surprising that the model achieves respectable results in the Grid-Search task, considering it must compute a vast number of grid points, which increase exponentially as the data per sample expands. The model faces challenges in the Hill-Climbing task, evident from a policy metric significantly exceeding zero. This suggests that meta-heuristics may pose greater difficulty for LLMs compared to other tasks.
Remarkably, GPT-turbo-3.5 and GPT-4 excel at performing Black-Box tasks in small-size samples. From Figure 3, we can see that GPT-turbo-3.5 performs notably with three data points, whereas GPT-4 excels at three and six data points. Interestingly, as the number of data points increases, the performance of both models gradually diminishes. Eventually, GPT-4 establishes itself as the superior performer.
Figure 3: GPTs average performance of Black-Box Task
@@
In the context of the Gradient-Descent Task, both models exhibit strong performance, with GPT-4 edging out GPT-turbo-3.5 by a slight margin. Figure 4 underscores this by revealing a policy metric that consistently hovers near zero, signifying a remarkable alignment between the LLM’s output and the ground truth. Despite a decline in the goal metric as the sample size increases, the consistently low and stable value of the policy metric underscores the fact that GPT’s performance in the gradient-descent task is nearly on par with the truth model or algorithm.
Figure 4: GPTs average performance of Gradient-Descent Task
6-2 Analysis
LLMs showcase optimization prowess across diverse domains. While not all solutions consistently achieve substantial improvements or align perfectly with ground truth, the predominantly positive goal metric values across most tasks and datasets indicate their capability for optimization. Among these tasks, Gradient Descent emerges as the leading performer, while the Hill-Climbing task poses greater challenges. Remarkably, GPT-4 even surpasses ground truth in a dataset comprising six data points, underscoring its exceptional aptitude for calculus. This underscores the notion that LLMs possess a versatile capacity to optimize within various problem spaces, which allows for the possibility of switching between different optimization methods within a single task.
LLM can be real optimizers. In the Black-Box Optimization task, LLMs are treated as black-box optimizers, which try to fit an unknown loss function. In each iteration, they are asked to offer their preferred solutions based on the discovered solutions and their true scores. Notably, when there are three data points, a more pronounced optimization capability is observed. Despite the performance decline with an increase in data points (variables), this observation intriguingly suggests that LLMs may possess inherent optimization capabilities of their own.
Data quantity influence stability. In cases where samples have fewer data points, it’s common to observe high uncertainty metric values accompanied by significant variations in policy and goal metrics. This suggests that while Large Language Models generally perform well in most trials, we may still encounter instances of failure, such as getting stuck in neither global nor local optima, optimizing in the wrong direction, and regenerating solutions that already appeared when attempting the same task again with the same sample. Figure 2 and 5 both highlight the conspicuous pattern of uncertainty, where the uncertainty initially rises and then gradually decreases. This behavior is consistent across different tasks and models, warranting further investigation to fully understand its underlying causes.
Figure 5: Uncertainty Metric of different tasks and models
Prompting methods may contribute to stability. In the Gradient-Descent task, we employ the self-consistency technique [4], where we conduct five repetitions for each iteration and select the solution that emerges most frequently. From Figure 6, we can see that GPT-4 performances increase largely, and the confidence interval for both the policy-metric and goal-metric narrows, indicating improved stability and reliability. Nonetheless, this approach does not yield favorable outcomes when applied to GPT-turbo-3.5. This suggests the need for further investigation within the realm of variance reduction to better understand the contributing factors to these results.
Figure 6: Average Performance of Gradient-Descent Task with Self-Consistency

### v3

@@
Abstract.
In this study, we evaluate the optimization capabilities of Large Language Models (LLMs) across diverse mathematical and combinatorial optimization tasks, where each task is described in natural language. These tasks require LLM to iteratively generate and evaluate solutions through interactive prompting, where each optimization step involves generating new solutions based on past results and then pass to subsequent iterations. We demonstrate that LLMs can perform various optimization algorithms and act as effective black-box optimizers, capable of intelligently optimizing unknown functions. We also introduce three simple yet informative metrics to evaluate optimization performance, applicable across diverse tasks and less sensitive to test sample variations. Our findings reveal that LLMs excel at optimizing small-scale problems with limited data and their performance is significantly affected by the dimension of problem and values, highlighting the need for further research in LLM optimization.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06copyright: none
1. Introduction
@@
Large Language Models have demonstrated exceptional capabilities in reasoning across a variety of natural language-based tasks  (Wei et al., 2022). However, their potential extends beyond multiple-choice questions or single-question answering. This work explores LLMs’ effectiveness in optimization across diverse tasks and problem dimensions. Optimization involves iteratively generating and evaluating solutions to improve a given objective function. Our research assesses LLM performance in interactive optimization, where each step generates new solutions based on previous ones and their values.
@@
We conduct our study with four different types of optimization algorithms: Gradient Descent, Hill Climbing, Grid Search, and Black Box Optimization. To provide a comprehensive evaluation of LLM performance, we introduce three distinct metrics. These metrics provide a multifaceted view of task performance and are applicable across a broad spectrum of optimization tasks, reducing sensitivity to sample variations.
Our findings suggest that LLMs show impressive optimization capabilities, especially in small-scale problems. However, their performance is notably affected by factors like sample size and value range. These observations underscore the need for further research within the domain of optimization tasks tailored for LLMs. It’s important to note that our work does not aim to outperform state-of-the-art optimization algorithms for either mathematical optimization or combinatorial optimization problems. Instead, our goal is to showcase the potential of LLM in these optimization domains and find out limitations in these settings.
Our contributions are summarized as follows:
  • Exploring the potential of LLMs in mathematical and combinatorial optimization scenarios.
  • Introduce three novel metrics for assessing LLM performance in optimization tasks.
  • Delve into factors that influence LLM performance using our metrics, with a particular emphasis on the impact of problem dimension and task type.
The remainder of this paper is structured as follows. In Section 2, we present preliminary works on LLMs for addressing optimization challenges. In Section 3, we defined 4 optimization algorithms in the case studies. In Section 4, we demonstrate that LLMs with iterative prompting strategy function as optimizers. In Section 5, we present three metrics that we have designed to assess the overall performance of LLMs in undertaking optimization tasks. Section 6, details our experimental results, showcasing the effectiveness of using LLMs as optimizers. In Section 7, we consolidated noteworthy observations and points of discussion from the experiments. Finally, Section 8 summarizes and concludes the paper.
2. Related Works
@@
In various optimization scenarios, the utilization of Large Language Models (LLMs) has become indispensable for the development of optimization algorithms or agent systems capable of handling complex and informative text-based feedback. In this section, we summarize three significant related works that leverage LLMs to tackle optimization and reinforcement learning challenges. These works showcase the adaptability and effectiveness of LLMs in addressing optimization and learning challenges across various domains.
@@
Optimization by PROmpting (OPRO) (Yang et al., 2023) OPRO harnesses LLMs as versatile optimizers by describing optimization tasks in natural language prompts. It iteratively generates and evaluates solutions from these prompts, demonstrating superior performance on tasks like linear regression and traveling salesman problems. OPRO outperforms human-designed prompts by up to 50% on challenging tasks.
@@
Reflexion (Shinn et al., 2023) Reflexion introduces a novel framework for training language agents that rely on linguistic feedback rather than traditional reinforcement learning. This framework delivers outstanding results, boasting a remarkable 91% pass@1 accuracy on coding tasks—an exceptional 11% improvement over previous state-of-the-art models. Reflexion’s success underscores the potential of linguistic feedback as a powerful training mechanism.
@@
EvoPrompt (Guo et al., 2023) EvoPrompt automates prompt optimization by connecting LLMs with evolutionary algorithms. This automated process surpasses human-designed prompts by up to 25% and outperforms existing automatic prompt generation methods by an impressive 14%. EvoPrompt’s success highlights the relationship between Large Language Models and traditional algorithms, showcasing the potential for enhanced problem-solving capabilities through this synergistic fusion.
3. Problem setting
@@
We design four optimization tasks that require the model to algorithmically search for the optimal value of parameters. These tasks encompass Gradient-Descent, Hill-Climbing, Grid-Search, and Black-Box Optimization, each representing unique optimization domains: gradient-based, meta-heuristics, decision-theoretic, and Bayesian. In terms of parameter types, Grid-Search and Hill-Climbing involve discrete search spaces, while Gradient-Descent and Black-Box Optimization tackle continuous search spaces. Following is detailed information on each optimization task.
@@
Gradient-Descent assesses the model’s proficiency in advanced calculations and its grasp of the principles of gradient descent. We instruct LLMs to undertake a conventional gradient descent optimization process based on the loss function they have defined. LLMs need to compute the gradient and update the parameters using the gradient information and the learning rate given.
@@
Hill-Climbing evaluate the LLM’s capability to adhere to custom predefined rules they have not seen before. LLMs start with an initial solution and iteratively explore nearby solutions by making small incremental changes. In our task, neighboring solutions are generated by selecting a specific element within the solution and either increasing or decreasing it by one each time. Subsequently, the neighbor solution with the minimum loss is chosen as the new solution and passed to the next iteration.
@@
Grid-Search assesses the LLM’s ability to conduct exhaustive searches and locate optimal solutions within a predefined search space. LLMs are tasked with generating all grid points and systematically searching for the point that results in the lowest loss according to the given loss function.
@@
Black-Box Optimization evaluates the LLM’s ability to make informed decisions and optimize in an abstract problem-solving context. We treat the LLMs as black boxes that try to fit an unknown loss function. We provide the LLM with a limited set of solutions, each paired with its respective true loss value. The LLM’s objective is to discover new solutions that have lower losses than the existing solutions in each iteration by themselves.
4. Methodologies
In this section, we show how LLMs, guided by iterative prompting, can effectively function as optimizers, akin to various optimization algorithms. To systematically navigate the search space, we introduce an iterative prompting framework that enables LLMs to incrementally achieve better solutions within the search space through iterative processes.
@@ 4. Methodologies
We applied Chain of Thoughts and iterative prompting as our prompting method. LLM will accomplish each step with reasoning thoughts as intermediate outputs. In each of these tasks (optimization algorithm), LLMs are initially required to formulate the loss function based on given samples. Then each optimization iteration is composed of two steps: (1) Generates new solution based on algorithm instructions and past search results (2) Calculate loss of new solution and add the results to the prompt of the next iteration. We keep repeating the two steps until the stop criteria are met. Figure 1 shows an overview of how LLM performs optimization in interactive settings.
@@
To create an interactive environment, we utilize the chat mode of GPTs, where the entire conversation history serves as the prompt. This allows LLMs to retain memory of past search results and reasoning paths. New instructions are appended to ongoing conversation records with each iteration. If the dialogue surpasses the token limit, earlier portions are removed.
Figure 1. Overview of our prompting strategy. (1) LLMs formulate the loss function based on given samples. (2) Given algorithm instructions and past results, LLM generates a new solution. (3) Calculate the loss of the new solution and add the solution-score pairs to the prompt of the next iteration. (4) Repeat the second and third steps until stop criteria are met.
5. Evaluation
We devised three novel metrics for the comprehensive evaluation of LLM capabilities. In this section, we will explain the design and objective of each metric. These metrics offer versatility in assessing LLM performance across diverse tasks, making concurrent evaluation easier. Their reliance on ratio measures, rather than differences, makes them less sensitive to sample variations.
5.1. Goal Metric
Goal metric evaluates how effectively LLMs perform optimization. It provides a quantitative measure of the degree to which the LLM contributes to minimizing the loss function values. In other words, ensuring that the ultimate solution loss is lower than the initial solution. We define the $goal$ $metric$ of a test sample $j$ as :
(1) $G_{j}=\frac{1}{N}\sum_{i=1}^{N}\frac{loss_{LLM,init}-loss_{LLM,i}}{loss_{LLM,init}}$
@@ 5.1. Goal Metric
5.2. Policy Metric
Policy metric assesses the degree of alignment between the final model output and the ground truth. Beyond self-improvement, which is measured by $goal$ $metric$, it is also crucial to appraise the LLMs’ capability to operate in a manner consistent with our truth model algorithm. This metric serves as an indicator of the LLM’s adeptness in adhering to task-specific instructions. We define the $policy$ $metric$ of a test sample $j$ as :
(2) $P_{j}=\frac{1}{N}\sum_{i=1}^{N}\frac{loss_{LLM,i}-loss_{truth}}{loss_{truth}}$
where $loss_{LLM,i}$ is the LLM output loss of trial $i$, $loss_{truth}$ is the ground truth of sample $j$ and $N$ is the number of trials. Since the policy metric measures the disparity between the ground truth and the LLM’s output, a lower policy metric value indicates a more effective alignment of the LLM’s actions with the prescribed guidelines. When the value is negative, it means that LLM’s performance surpasses the ground truth.
5.3. Uncertainty Metric
Uncertainty metric quantifies the variability in the LLM’s solutions under identical conditions. Stability is a crucial characteristic in optimization tasks. We hope that the LLMs produce identical results in every trial involving the same sample, even under conditions with temperatures greater than zero. We define the $uncertainty$ $metric$ of a test sample $j$ as :
(3) $U_{j}=\frac{1}{N}\sum_{i=1}^{N}(loss_{LLM,i}-\overline{loss_{LLM}})^{2}$
@@ 5.3. Uncertainty Metric
6. Experiments
This section provides details of our experimental configurations and highlights the outcomes of experiments. Subsection  6.1 outlines the process of generating synthetic datasets for all optimization tasks, while subsection  6.2 elucidates the detailed settings of our experiment. Lastly, subsection  6.3 offers a concise summary of the outcomes derived from our experiment.
6.1. Dataset
In the experiment, we create five datasets with $d$ values chosen from the set $\{3,6,12,24,48\}$ and generate instances belonging to $[0,10]^{d}$ in each dataset to examine sensitivity to the number of parameters, representing the dimension of the optimization problem. For instance, $d=3$ indicates that there are 3 variables in the loss function and the dimension of this optimization problem is 3. We then apply each instance to a loss function and find the true solution for each parameter search task. These authenticated solutions, coupled with their associated losses, not only serve as the ground truth for the tasks but also act as a pivotal benchmark against which the solutions derived by LLMs are systematically evaluated and compared in the ensuing analysis.
6.2. Detailed Settings
In our experiment, We set the LLM temperature to 0.8 and the reset as default. We performed 5 repetitions of the test for each instance in the dataset, with the LLM conducting 10 iterations of parameter search in each repetition. We excluded excessively biased results to prevent our metrics from being skewed by a minority of poorly performing test outcomes. All experiments employ the GPT-turbo-3.5 ’0613’ version as the Language Model.
6.3. Main Results
@@
We summarize the outcomes of our experiment and subsequently examine the common trends observed across all experiments. In every plot, the x-axis displays the dimension of the optimization problem. In the case of the goal metric and policy metric plots, the y-axis illustrates the average metric value for the respective tasks, while the shaded area in a lighter color delineates the confidence interval of the metric, denoted as $[value-std,value+std]$. As for the uncertainty metric plot, the y-axis showcases the uncertainty metric value, which corresponds to the standard deviation of the LLM final solution loss. It is worth noting that the Goal Metric graph excludes the non-iterative Grid-Search task due to its non-iterative nature, while the Policy Metric graph omits the Black-Box task due to unattainable ground truth.
@@
LLMs show strong optimization capabilities in small-scale problems. Our experiments test the comprehensive optimization capabilities of LLMs. Observing figure 2, GPT-turbo-3.5 showcases considerable optimization capabilities across various scenarios. Impressively, in the Gradient-Descent task, GPT-turbo-3.5 even surpasses the ground truth, particularly in the case of the sample dimension equal to six. It’s also surprising that the model achieves respectable results in the Grid-Search task, considering it must compute a vast number of grid points, which increase exponentially as the dimension of the problem expands. The model faces challenges in the Hill-Climbing task, evident from a policy metric significantly exceeding zero. This suggests that meta-heuristics may pose greater difficulty for LLMs compared to other tasks.
Figure 2. Goal Metric and Policy Metric hover from positive to near zero, signifying substantial optimization capability and alignment between LLM’s output and ground truth.
LLMs show potential as Black-Box Optimizers. Favorable performance in Black-Box experiments suggests the use of LLM as an optimizer without giving any algorithm instructions. From figure 3, we can see that GPT-turbo-3.5 performs notably when the dimension of the problem is three, whereas GPT-4 excels when the dimensions are three and six. Interestingly, as the dimension increases, the performance of both models gradually diminishes. Eventually, GPT-4 edged out GPT-turbo-3.5 by a slight margin in optimization and stability.
Figure 3. Goal Metric reflects the performance of LLMs as Black-Box optimizer, showing strong performance with instances of smaller dimensions.
@@
LLMs exhibit strong performance in Gradient-Descent. Gradient-Descent experiment tests the model’s proficiency in advanced calculations and grasp of mathematics principles. Figure 4 underscores this by revealing a policy metric that consistently hovers near zero, signifying a remarkable alignment between the LLM’s output and the ground truth. Despite a decline in the goal metric as the sample size increases, the consistently low and stable value of the policy metric underscores the fact that GPT’s performance in the gradient-descent task is nearly on par with the truth model.
Figure 4. Low values in the Policy Metric and high positive values in the Goal Metric indicate the robust performance of the LLM in the gradient descent task.
7. Analysis and Discussion
In this section, we consolidate several crucial insights derived from our experimental results and subject them to analysis.
Pretrained Knowledge dominates the optimization capability of LLM. Among all optimization tasks performed by LLMs, Gradient Descent emerges as the leading performer, while Hill-Climbing poses greater challenges. The main difference between the two tasks is that Hill-Climbing is a heuristic algorithm with more user-specific parameters, whereas gradient descent is an optimization algorithm that relies more on mathematical principles. This suggests that LLM optimization capabilities primarily stem from pretrained knowledge stored within the model parameters, rather than from context knowledge provided by users. Our findings align with previous research  (Chen et al., 2022; Pagnoni et al., 2021; Zhou et al., 2023) showing that language models often prioritize their prior knowledge over new context. Achieving balanced attention to both prior and context knowledge is essential for further research to improve the optimization capability of language models.
LLMs are potential hybrid optimizers. The predominantly positive goal metric values across most tasks and datasets indicate LLMs’ capability for optimization. This highlights their versatile capacity to optimize across different problem spaces, potentially allowing for the switching between optimization methods within a single task. Such switching can help LLMs better explore the solution space and escape local optima where they might get stuck. This is a significant advantage of LLMs in optimization, as they can easily change methods through a simple natural language prompt during iterations. Furthermore, LLMs can act as agents (world models) that use different algorithms as tools (actions), switching methods by evaluating the optimization path from past to present (state). This adaptability underscores the potential of LLMs to enhance optimization processes through dynamic method selection and strategic problem-solving.
LLMs possess richer solution space in small-scale problems. In our experiments, we observed high uncertainty metric values and significant variations in policy and goal metrics when samples had smaller dimensions. Interestingly, LLMs tend to perform more effectively with smaller dimension instances, suggesting a correlation between higher uncertainty and better performance. This consistent pattern across various tasks and models indicates that LLMs have a richer solution space when tackling small-scale problems. The expanded solution space leads to higher uncertainty, providing LLMs with a broader range of solutions to explore. This highlights the importance of dimension reduction in data preprocessing for effective optimization by LLMs. Figure 2 and 5 both highlight the pattern of uncertainty, where the uncertainty initially rises and then gradually decreases.
Figure 5. An initial rise followed by a decline in the Uncertainty Metric with instance dimension growth suggests LLMs may have a richer sample space for small-scale problems, consistent across tasks and models.